# Dynamics of Agentic Loops in Large Language Models: A Geometric Theory of Trajectories From Semantic Contraction to Exploratory Divergence

### v3

@@ 5.3 Metrics and Methodology
We also compute the calibrated similarity:
$\tilde{s}_{i}=\tilde{s}(e_{i},e_{i}^{\prime})=f_{\text{isotonic}}(s_{i}^{(m)}),$ (14)
where $f_{\text{isotonic}}$ denotes the isotonic regression mapping introduced in Section 4.
For each perturbation type $k$ with $N_{k}$ pairs, we compute three stability indicators:
$\mu_{k}=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}s_{i},\quad\sigma_{k}=\sqrt{\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}(s_{i}-\mu_{k})^{2}},\quad r_{k}=\frac{1}{N_{k}}\sum_{i\in\mathcal{I}_{k}}\mathbf{1}_{\{s_{i}\geq\tau\}},$ (15)
where $\mathcal{I}_{k}$ denotes the index set of pairs for perturbation type $k$ and $\tau$ is the stability threshold.
Threshold selection.
Rather than choosing arbitrary thresholds, we use the data-driven high-confidence similarity thresholds defined in Section 3.6:
  • For base model similarity: $\tau=\tau_{\text{HCS}}^{(m)}\approx 0.72$
  • For calibrated similarity: $\tau=\tilde{\tau}_{\text{HCS}}\approx 0.65$
The stability rate $r_{k}$ thus measures the proportion of perturbed pairs that maintain similarity above the threshold derived from high-confidence human judgments, providing a principled criterion for local stability.
5.4 Empirical Results
Tables 2 and 3 report the local stability metrics for base model (raw cosine) and calibrated similarities, respectively. For each perturbation type, we provide the number of pairs, mean similarity, standard deviation and stability rate (proportion of pairs exceeding the corresponding threshold).
Perturbation Type Pairs ($N_{i}$) Mean ($\mu_{\text{base},i}$) Std. Dev. ($\sigma_{\text{base},i}$) Stab. Rate ($r_{\text{base},i}$)
DETERMINER VARIATION 100 0.932 0.017 1.00
TENSE VARIATION 100 0.909 0.030 1.00
SYNONYM SUBSTITUTION 100 0.853 0.060 0.95
LOGICAL PARAPHRASE 100 0.891 0.023 1.00
NOMINALIZATION 100 0.866 0.035 1.00
COREFERENCE EXPANSION 100 0.882 0.022 1.00
QUANTIFIER VARIATION 100 0.848 0.039 0.97
All types (overall) 700 0.883 0.045 0.99
Table 2: Local stability metrics using base model similarity $s^{(m)}$ (threshold $\tau_{\text{HCS}}^{(m)}=0.72$).
Perturbation Type Pairs ($N_{i}$) Mean ($\mu_{\text{sim},i}$) Std. Dev. ($\sigma_{\text{sim},i}$) Stab. Rate ($r_{\text{sim},i}$)
DETERMINER VARIATION 100 0.926 0.054 1.00
TENSE VARIATION 100 0.868 0.069 1.00
SYNONYM SUBSTITUTION 100 0.799 0.082 0.87
LOGICAL PARAPHRASE 100 0.831 0.049 1.00
NOMINALIZATION 100 0.818 0.078 1.00
COREFERENCE EXPANSION 100 0.830 0.045 1.00
QUANTIFIER VARIATION 100 0.798 0.050 0.97
All types (overall) 700 0.839 0.075 0.98
Table 3: Local stability metrics using calibrated similarity $\tilde{s}$ (threshold $\tilde{\tau}_{\text{HCS}}=0.65$).
5.5 Discussion
The comparison between Tables 2 and 3 reveals two critical findings regarding local stability.
Embedding function exhibits strong local stability.
The base model similarity analysis (Table 2) demonstrates that the embedding function $\psi$ exhibits excellent local stability, with an overall stability rate of $99\%$. All perturbation types maintain mean similarities well above the threshold $\tau_{\text{HCS}}^{(m)}=0.72$, ranging from $0.848$ (QUANTIFIER_VARIATION) to $0.932$ (DETERMINER_VARIATION). This confirms that the pretrained embedding model is robust to small linguistic perturbations, preserving high semantic similarity even under controlled transformations such as tense variation, synonym substitution and logical paraphrasing.
Isotonic calibration introduces minimal stability degradation.
Table 3 shows that isotonic regression calibration slightly reduces the overall stability rate from $99\%$ to $98\%$, a modest degradation of only 1 percentage point. This small decrease is attributable to the piecewise constant (discontinuous) nature of isotonic regression: when a perturbed pair $(a_{i},a_{i}^{\prime})$ has a raw cosine similarity $s^{(m)}$ near a discontinuity point of $f_{\text{isotonic}}$, the calibrated similarities $\tilde{s}(e_{i},e_{i})$ and $\tilde{s}(e_{i}^{\prime},e_{i}^{\prime})$ may fall into different constant segments, potentially crossing the threshold $\tilde{\tau}_{\text{HCS}}=0.65$.
The most affected perturbation type is SYNONYM_SUBSTITUTION, whose stability rate drops from $0.95$ to $0.87$. This is expected: synonym substitutions produce moderate cosine similarities (mean $0.853$) that fall in a region where isotonic regression has more step discontinuities. In this moderate range, small linguistic perturbations can cause similarities to jump across step boundaries in $f_{\text{isotonic}}$, leading to larger changes in calibrated values than in the original cosine similarities. Conversely, perturbation types with very high base similarities (TENSE_VARIATION: $0.909$, DETERMINER_VARIATION: $0.932$) remain perfectly stable ($r=1.00$) after calibration, as their similarities fall in a region where isotonic regression is nearly constant and far from the threshold.
Practical acceptability of calibrated stability.
Despite the theoretical discontinuity of isotonic regression, the empirical stability rate of $98\%$ remains highly acceptable for practical use. This demonstrates that isotonic regression achieves its primary goal (eliminating bias and improving correlation with human judgments) without substantially compromising local stability. The 1% degradation is a negligible cost compared to the significant gains in calibration quality (ECE reduction from $0.0797$ to $0.0000$, as shown in Table 1).
Implications for dynamic analysis.
These results validate the use of calibrated similarity $\tilde{s}$ for studying agentic loop dynamics. The embedding function $\psi$ exhibits strong local stability and isotonic calibration preserves this property with only minimal degradation. This ensures that the semantic trajectories observed in agentic loops reflect genuine semantic evolution rather than instabilities introduced by either the embedding function or the calibration process. With local stability empirically established, we now turn to the formal analysis of agentic loop dynamics, confident that our similarity metric provides a stable and human-aligned foundation for measuring iterative transformations in generative systems.
@@ 6 Agentic Loop Dynamics: Theoretical Framework
6.1 Foundational Spaces and Operations
We begin by establishing the mathematical spaces and primitive operations that underlie all agentic loop structures.
Artifact and embedding spaces.
Let $\mathcal{A}$ denote the artifact space, where each element $a\in\mathcal{A}$ represents a linguistic artifact (e.g., a sentence, paragraph or document). Let $\psi:\mathcal{A}\to\mathcal{E}$ be the embedding map introduced in Section 3, which projects artifacts into a semantic vector space where their meaning can be quantitatively compared.
Primitive operations.
We introduce two fundamental operations:
@@ 7.2.1 Global and Local Dynamics
Figure 9 presents the evolution of both local and global geometric metrics along the trajectory of 50 iterations. These indicators provide a direct view of the loop’s dynamics before any segmentation into clusters.
Figure 9: Local and global geometric evolution of the contractive loop trajectory. (Top) Local distances and similarities between consecutive embeddings $(e_{t},e_{t-1})$. (Bottom) Global distances and similarities relative to the initial embedding $e_{0}$.
@@
The local Euclidean distance $\|e_{t}-e_{t-1}\|_{2}$ fluctuates between $0.45$ and $0.55$, while the corresponding calibrated similarity $\tilde{s}(e_{t},e_{t-1})$ remains consistently high in the range 0.820.820.950.95. This indicates that each rewriting step introduces only minimal changes to the representation, consistent with the model’s smooth behavior under incremental rewriting.
@@
When compared to the initial state $e_{0}$, the global distance stabilizes around $0.6$ and the calibrated similarity remains near $0.75$. After a few initial iterations, the trajectory stops drifting and remains confined within a bounded region of $\mathcal{E}$, consistent with the expected contractive regime. This global stability validates the use of local geometric measures as reliable indicators of semantic convergence.
@@ 7.3.1 Global and Local Dynamics
Figure 13 presents the evolution of both local and global geometric metrics over 50 iterations. Unlike the contractive regime, these indicators show strong irregularity and the absence of any stabilization phase.
Figure 13: Local and global geometric evolution of the exploratory loop trajectory. (Top) Local distances and similarities between consecutive embeddings $(e_{t},e_{t-1})$. (Bottom) Global distances and similarities relative to the initial embedding $e_{0}$.
@@
The local Euclidean distance $\|e_{t}-e_{t-1}\|_{2}$ remains high across iterations, oscillating between $0.8$ and $1.3$ with no decreasing trend. Similarly, the local similarity $\tilde{s}(e_{t},e_{t-1})$ fluctuates widely between $0.2$ and $0.6$, occasionally dropping close to zero. These erratic variations indicate that each step produces a major semantic reorientation rather than a refinement of the previous state.
@@
The global distance to the initial embedding $e_{0}$ rapidly increases during the first iterations and stabilizes around $\|e_{t}-e_{0}\|_{2}\approx 1.3$ (see Figure 13, bottom left). The corresponding global similarity $\tilde{s}(e_{t},e_{0})$ remains very low (mostly below $0.2$), consistent with a loop that continuously generates artifacts semantically unrelated to the starting point. This behavior reveals the absence of any long-term memory or attractor basin in the semantic manifold.
@@ 7.3.2 Clustering Attempt and Absence of Attractors
The exploratory loop displays:
  • persistent high local distances ($\|e_{t}-e_{t-1}\|_{2}\approx 1.0$) and low similarities ($\tilde{s}<0.5$);
  • no geometric stabilization, either locally or globally;
  • and the complete absence of clusters or attractor basins for all tested parameters.
In semantic terms, this loop corresponds to a continuous destruction and regeneration of meaning: each iteration rewrites the narrative space from scratch, producing an intrinsically non-convergent agentic dynamic.
@@ 7.4 Comparative Summary
The empirical analysis of contractive and exploratory loops reveals fundamentally distinct dynamical regimes, following the formal definitions established in Section 6.5. Table 9 summarizes the key quantitative and qualitative differences between these two archetypal behaviors.
@@
Mean local distance $\sim$0.5 $\sim$1.0
Mean local similarity >>0.85 <<0.5
@@
Global drift Bounded Unbounded
Regime type Convergent Divergent
@@ 8.3 Toward Geometry of Agentic Cognition
8.4 Application: Creativity as Objective
The agentic geometric cognition framework introduced in Section 8.3 raises a natural application question: what if the objective that the agent targets is to maximize creativity? This section explores how computational creativity, formalized by Boden Boden (2009) as exploration-transformation-combination processes and by Colton and Wiggins Colton and Wiggins (2012) as novelty-value-surprise dimensions, can be operationalized as a specific objective within the geometry of agentic cognition framework.
Creativity metrics as geometric observables.
To target creativity as an objective, an agent with geometric self-awareness must first be able to measure creativity dimensions along its trajectory. Our geometric framework provides measurable indicators that correspond to Colton & Wiggins’ creativity dimensions:
  • Novelty corresponds to exploratory dynamics, that is, trajectories that visit diverse regions of semantic space without settling into stable attractors. The absence of persistent clusters and low calibrated similarity between consecutive states ($\tilde{s}(e_{t},e_{t-1})$) indicate ongoing exploration that generates new ideas rather than refining existing ones.
  • Value corresponds to convergence toward stable attractors with low dispersion. Trajectories that stabilize around a fixed point produce semantically coherent, trustworthy outputs. Conversely, divergent trajectories exhibit semantic drift, where outputs lose coherence and become unreliable, yielding low value.
  • Surprise corresponds to sudden, unpredictable semantic transitions, specifically trajectory discontinuities where the semantic state jumps sharply after periods of stability. Contractive loops exhibit low surprise (smooth, predictable convergence), while exploratory loops may produce high surprise through unpredictable jumps between distant semantic regions.
This operationalization enables measurable creativity control. Boden’s exploratory creativity can be triggered within agentic loop iterations by introducing high-temperature sampling and role-based divergent prompts Zhao et al. (2024) (for instance, adopting perspectives such as an artist, a child or a poet), which induce high dispersion and novelty generation. However, unconstrained exploration risks semantic drift and loss of value. To balance creativity, subsequent iterations can act as consolidators, employing role-based contractive prompts (e.g., scientist, engineer, critic) that reduce dispersion and restore coherence, maximizing value while preserving novel content.
Composite loops as creativity engines.
This dual-phase architecture naturally maps to composite loops as defined in Section 6.2. A composite loop with two phases per iteration would take the form:
$a_{t+1}=\mathrm{LLM}_{\text{contract}}(P_{\text{contract}}(\mathrm{LLM}_{\text{explore}}(P_{\text{explore}}(a_{t})))),$ (27)
where $P_{\text{explore}}$ uses role-based divergent prompts (e.g., "adopt the perspective of an artist") to induce exploratory divergence and $P_{\text{contract}}$ employs precision-oriented prompts (e.g., "critically evaluate and consolidate") to enforce contraction. The generation parameters can be tuned accordingly: $\mathrm{LLM}_{\text{explore}}$ could use high temperature (e.g., $T=1.2$) for creative diversity, while $\mathrm{LLM}_{\text{contract}}$ could use low temperature (e.g., $T=0.3$) for deterministic refinement. The singular loop experiments presented in Section 7 establish that these individual $(P,\mathrm{LLM})$ pairs can reliably induce their respective regimes. Composing them sequentially would create an oscillatory dynamic where each iteration explores the semantic manifold then contracts toward coherent outputs, effectively implementing Boden’s explore-transform-combine cycle through geometric dynamics.
The interplay between these regimes may enable a new kind of controlled computational creativity, where the balance between exploration and consolidation is explicitly managed through prompt design, role assignment and iterative feedback. Future work could formalize this balance through composite loop architectures that adaptively modulate their dynamics based on creativity objectives, bridging Boden’s theoretical framework with the geometric formalism introduced here.
@@ 9 Conclusion
This work introduced a geometric framework for analyzing the dynamics of agentic loops: recursive processes in which large language models iteratively transform their own outputs. By grounding these dynamics in a calibrated semantic embedding space (,s~)($\mathcal{E}$$\tilde${s}) that enables human-aligned measurements, we made it possible to quantify how iterative LLM processes unfold, stabilize or diverge across iterations.
@@
    A formal distinction between the artifact space $(\mathcal{A},F)$, where linguistic transformations occur, and its continuous representation space $(\mathcal{E},\|\cdot\|)$, enabling quantitative measurement of agentic dynamics through geometric trajectory analysis.
@@
    A calibrated similarity metric $\tilde{s}=f_{\mathrm{isotonic}}\circ s^{(m)}$ that corrects the anisotropy of embedding spaces and provides human-aligned geometric indicators: dispersion, displacement, cluster coherence and trajectory discontinuities.
@@
    An empirical demonstration that agentic loops exhibit distinct dynamical regimes (contractive or exploratory) depending solely on their prompt formulation, with measurable geometric signatures that characterize each regime.
@@
Using a controlled experimental setup, we showed that:
  • The contractive loop, driven by incremental rewriting, produces a smooth and bounded trajectory in $\mathcal{E}$ that converges toward a dense attractor region. Its dynamics exhibit local contractivity: successive embeddings remain close, dispersion decreases over time, and the trajectory stabilizes within persistent clusters.
  • The exploratory loop, built on alternating summarization and negation, generates an unbounded trajectory with no persistent clusters or attractors. Its behavior is characterized by high local displacements and low inter-iteration similarity, with dispersion remaining elevated throughout the process.
Together, these two regimes define a spectrum of agentic dynamics, from convergent stabilization to divergent exploration, each with distinct geometric signatures.
@@
From a dynamical systems perspective, these observations are consistent with a view of LLM-based transformations $F$ as mappings whose dynamics resemble those governed by an effective contraction rate:
$\|F(a_{1})-F(a_{2})\|\leq L\,\|a_{1}-a_{2}\|,\qquad L<1\Rightarrow\text{contraction},\quad L>1\Rightarrow\text{divergence}.$ (28)
Since $F$ operates in $\mathcal{A}$ but all geometric measurements occur in $\mathcal{E}$ through $\psi$, the contraction rate is understood with respect to the induced operator $\hat{F}=\psi\circ F$. While such bounds cannot be measured directly for deep generative operators, the observed trajectory patterns suggest a regime-dependent effective contraction rate that could be estimated through embedding-space analysis. This opens the door to a formal theory of agentic dynamics, linking prompt design, model sampling and geometric regularity.
@@
The contrast between contractive and exploratory regimes highlights a key tension in autonomous generative systems: the need to balance semantic exploration and stability. While this paper focused on singular loops to establish foundational dynamical phenomena, the natural next step is the study of composite loops, structures where multiple LLM calls with different prompts compose sequentially within each iteration:
$a_{t+1}=\mathrm{LLM}_{2}(P_{2}(\mathrm{LLM}_{1}(P_{1}(a_{t})))).$ (29)
For example, a generate-critique-revise cycle would use $P_{1}$ for generation, $P_{2}$ for critique and $P_{3}$ for revision. The singular loop experiments demonstrated that individual $(P,\mathrm{LLM})$ pairs can reliably induce contractive or exploratory dynamics. Composing these building blocks offers a systematic path toward controlled agentic behavior: alternating exploratory (divergent) and consolidating (contractive) phases to enable agents that explore the semantic space while maintaining coherence.
This principle extends naturally to complex agentic architectures, where multiple loops interact, synchronize or self-regulate within shared semantic manifolds. Hierarchical, cooperative and adversarial multi-loop systems could exhibit emergent behaviors (collective attractors, dynamic equilibria or oscillatory states) that go beyond the capabilities of individual loops. Understanding these phenomena requires extending the geometric framework presented here to handle compositions, feedback and synchronization across interacting semantic trajectories.
Final remark.
Ultimately, this work provides both a language and a methodology for describing the internal geometry of agentic loops in LLMs. By establishing these foundations through singular loop analysis, it opens the path toward understanding composite and complex agentic systems that balance exploration and consolidation. The framework connects naturally to computational creativity research, where the geometric indicators developed here (dispersion, cluster coherence, trajectory discontinuities) can operationalize theoretical notions of novelty, value and surprise. More broadly, it establishes a bridge between agentic AI and dynamical systems theory, encouraging future study of how autonomous generative processes evolve within semantic embedding spaces.

### v4

@@ 3 Semantic Space and Calibrated Similarity
Raw cosine similarity between embeddings suffers from anisotropy: pretrained embedding spaces concentrate vectors in a narrow cone, causing similarities to cluster around high values even for semantically distant pairs Ethayarajh (2019). To address this, we employ the calibrated similarity function proposed in Tacheny (2025):
$\tilde{s}(e_{1},e_{2})=f_{\mathrm{isotonic}}(\langle e_{1},e_{2}\rangle),$ (1)
where $f_{\mathrm{isotonic}}$ is a monotonic mapping learned via isotonic regression on human-annotated sentence pairs. This calibration eliminates systematic bias, improves correlation with human judgments (Spearman $\rho\approx 0.86$) and preserves local stability under small linguistic perturbations ($\geq 98\%$). Throughout this paper, all geometric measurements use the calibrated similarity $\tilde{s}$.
High-confidence similarity threshold.
Following Tacheny (2025), we use the data-driven decision boundary $\tilde{\tau}_{\text{HCS}}$, the 5% quantile of calibrated similarity among pairs with high human similarity. This threshold guarantees that at least 95% of truly similar pairs exceed this value and is used for cluster detection in subsequent sections. Note that the specific value of $\tilde{\tau}_{\text{HCS}}$ depends on the embedding model; throughout this paper, we adopt the same embedding model as Tacheny (2025), yielding $\tilde{\tau}_{\text{HCS}}\approx 0.65$.
Robustness to metric choice.
While we use calibrated similarity for improved interpretability, the qualitative conclusions of this work—contractive vs. exploratory dynamics, attractor formation, regime classification—depend only on order preservation, not on the specific numeric values. Since isotonic calibration is a monotonic transformation, it preserves pairwise similarity rankings and nearest-neighbor relationships. Consequently, all definitions and empirical trends reported here would hold under raw cosine similarity or any other order-preserving metric. The calibration improves human interpretability of threshold values but is not required for the framework’s validity.
@@ 4 Agentic Loop Dynamics: Theoretical Framework
4.1 Primitive Operations
Building on the artifact space $\mathcal{A}$, representation space $\mathcal{E}$, and calibrated similarity $\tilde{s}$ defined in Section 3, we introduce two fundamental operations:
@@ 5.2.1 Global and Local Dynamics
The calibrated similarity $\tilde{s}(e_{t},e_{t-1})$ between consecutive embeddings remains consistently high, in the range 0.820.820.950.95 (Figure 1). Each rewriting step introduces only minimal semantic changes.
Figure 1: Local similarity $\tilde{s}(e_{t},e_{t-1})$ for the contractive loop. Values remain consistently high (>0.8>0.8), indicating small semantic displacements between consecutive iterations.
@@
The similarity to the initial embedding $\tilde{s}(e_{t},e_{0})$ stabilizes near $0.75$ after initial iterations (Figure 2). The trajectory stops drifting and remains confined within a bounded region of $\mathcal{E}$.
Figure 2: Global similarity $\tilde{s}(e_{t},e_{0})$ for the contractive loop. The trajectory stabilizes at moderate similarity to the initial state, indicating bounded drift.
@@ 5.3.1 Global and Local Dynamics
The local similarity $\tilde{s}(e_{t},e_{t-1})$ fluctuates widely between $0.2$ and $0.6$, occasionally dropping close to zero (Figure 6). Each step produces a major semantic reorientation rather than incremental refinement.
Figure 6: Local similarity $\tilde{s}(e_{t},e_{t-1})$ for the exploratory loop. Values fluctuate erratically, indicating large semantic displacements at each iteration.
@@
The similarity to the initial embedding $\tilde{s}(e_{t},e_{0})$ remains very low throughout, mostly below $0.2$ (Figure 7), indicating continuous generation of artifacts semantically unrelated to the starting point.
Figure 7: Global similarity $\tilde{s}(e_{t},e_{0})$ for the exploratory loop. The trajectory rapidly diverges from the initial state and remains semantically distant throughout.
@@ 5.3.2 Clustering Attempt and Absence of Attractors
The exploratory loop displays persistently low local similarities ($\tilde{s}(e_{t},e_{t-1})<0.5$), no geometric stabilization either locally or globally, and complete absence of clusters or attractor basins for all tested parameters. In semantic terms, this loop corresponds to continuous destruction and regeneration of meaning: each iteration rewrites the narrative space from scratch.
@@ 5.4 Comparative Summary
The empirical analysis of contractive and exploratory loops reveals fundamentally distinct dynamical regimes, following the formal definitions established in Section 4.5. Table 6 summarizes the key differences.
@@
Mean local similarity $\tilde{s}(e_{t},e_{t-1})$ >>0.85 <<0.5
Global similarity $\tilde{s}(e_{t},e_{0})$ $\sim$0.75 <<0.2
@@
Regime type Contractive Exploratory
@@ 6.3 Toward Geometry of Agentic Cognition
6.4 Speculative Application: Creativity
The geometric indicators developed here—drift, dispersion, attractor formation—bear a suggestive correspondence to dimensions of computational creativity as formalized by Boden Boden (2009) and Colton & Wiggins Colton and Wiggins (2012). Exploratory dynamics (high dispersion, no stable attractors) may relate to novelty generation, while contractive dynamics (convergence, low dispersion) may relate to value consolidation.
This suggests a speculative direction: composite loops that alternate exploratory and contractive phases could potentially balance novelty and coherence, with prompt design and temperature modulating the regime at each phase Zhao et al. (2024). We emphasize that this remains a hypothesis for future investigation; the present work provides only the measurement infrastructure, not empirical validation of creativity-specific claims.
@@ 7 Conclusion
This work introduced a geometric framework for analyzing the dynamics of agentic loops: recursive processes in which large language models iteratively transform their own outputs. By grounding these dynamics in a calibrated semantic embedding space (,s~)($\mathcal{E}$$\tilde${s}), we made it possible to quantify how iterative LLM processes unfold, stabilize or diverge.
@@
    A formal theoretical framework distinguishing artifact space $(\mathcal{A},F)$ from representation space $(\mathcal{E},\|\cdot\|)$, with precise definitions of trajectories, clusters, attractors and dynamical regimes.
@@
    An incremental cluster detection algorithm that operationalizes these concepts for empirical analysis.
@@
    Empirical demonstration that agentic loops exhibit distinct dynamical regimes (contractive or exploratory) depending on prompt formulation, with measurable geometric signatures.
@@
  • The contractive loop produces a bounded trajectory converging toward a dense attractor region, with local contractivity and persistent clustering.
  • The exploratory loop generates an unbounded trajectory with no persistent clusters, characterized by high local displacements and continuous semantic drift.
These regimes define a spectrum from convergent stabilization to divergent exploration.
@@
The observations are consistent with a view of LLM transformations as mappings whose dynamics resemble those governed by an effective contraction rate. While such bounds cannot be measured directly for deep generative operators, observed trajectory patterns suggest a regime-dependent effective contraction rate that could be estimated through embedding-space analysis.
@@
The contrast between contractive and exploratory regimes highlights a key tension: the need to balance semantic exploration and stability. Composite loops that alternate exploratory and contractive phases offer a systematic path toward controlled agentic behavior: exploring semantic space while maintaining coherence. This connects naturally to computational creativity research, where the geometric indicators developed here can operationalize theoretical notions of novelty, value and surprise.
More broadly, this work establishes a bridge between agentic AI and dynamical systems theory, encouraging future study of how autonomous generative processes evolve within semantic embedding spaces.